OpenStudy (faiithhh):

HOW WOULD I DO THIS IN THE GUESS METHOD? A motor is used to lift a person sitting on a platform that, combined weighs 650 N. The platform is lifted 25.0 m straight up in 18.0 seconds. What power is developed by the motor in watts?

OpenStudy (faiithhh):

i know that my equation will be p = w/t. i also know that t = 18. but i dont know what w is?

OpenStudy (faiithhh):

@sapphiremoon

OpenStudy (mathmate):

hint: w(work)=force \(\times\) distance joules (J). In the case of vertical lift (upwards), force = mg = weight in newtons (N) distance = 25 m (for this problem). Power = Work/time watts (W).

OpenStudy (osprey):

GUESS METHOD - never heard of it. Gauss, yes, but not guess. Except that it's what I do all the time.

OpenStudy (sapphiremoon):

Gosh, not familiar with work/power yet, I could try to go off what mathmate said, but you probably don't want me making stuff up as I go along XD. @faiithhh

OpenStudy (irishboy123):

7.4

OpenStudy (mathmate):

Hm... what unit would that be in?

OpenStudy (irishboy123):

Coulombs per inch would be my guess

OpenStudy (mathmate):

Ok, so you were using the Guess method! lol

OpenStudy (osprey):

@IrishBoy123 grams of EU cake per slug per parsec per metre squared ?